# Joint effective equidistribution of partial lattices in positive characteristic

### v2

@@ 1 Introduction
We endow $K_{\nu}^{\;\mathbf{d}}$ with the supremum norm, and any $K_{\nu}$-linear subspace of $K_{\nu}^{\;\mathbf{d}}$ with its induced norm and with its associated normalized Haar measure (giving mass one to its closed unit ball), see Subsection 2.2. A partial $R_{\nu}$-lattice $\Lambda$ of rank $\mathfrak{m}$} (or $\mathfrak{m}$}-lattice for short) in $K_{\nu}^{\;\mathbf{d}}$ is a discrete free $R_{\nu}$-submodule of rank $\mathfrak{m}$} generating a $\mathfrak{m}$}-dimensional $K_{\nu}$-vector subspace $V_{\Lambda}$ of $K_{\nu}^{\;\mathbf{d}}$. We denote by $\operatorname{Covol}(\Lambda)$ the covolume of $\Lambda$ in $V_{\Lambda}$. We say that $\Lambda$ is primitive if it is a free direct factor of $R_{\nu}^{\;\mathbf{d}}$. Among all the definitions of primitiveness that were equivalent for partial $\mathbb{Z}$}-lattices in ${\mathbb{R}}^{n}$ and non longer are, this turns out to be the appropriate one. We denote by $\operatorname{{\cal P}\!{\cal L}}_{{\mathfrak{m}},\mathbf{d}}$ the set of primitive $\mathfrak{m}$}-lattices in $K_{\nu}^{\;\mathbf{d}}$. For every point $y$ of every measurable space, we denote by $\Delta_{y}$ the unit Dirac mass at $y$.
@@
Let $\Lambda\in\operatorname{{\cal P}\!{\cal L}}_{{\mathfrak{m}},\mathbf{d}}$ be such that there exists $i\in{\mathbb{Z}}$ with $\frac{\operatorname{Covol}(\Lambda)}{\operatorname{Covol}(R_{\nu}^{\;{\mathfrak{m}}})}=q_{\nu}^{\;\operatorname{lcm}({\mathfrak{m}},{\mathfrak{n}})\,i}$ and $g\in\operatorname{GL}_{\mathbf{d}}({\cal O}_{\nu})$ sending $V_{\Lambda}$ to $K_{\nu}^{\;{\mathfrak{m}}}\times\{0\}$ (so that $g\Lambda$ becomes a full $R_{\nu}$-lattice of $K_{\nu}^{\;{\mathfrak{m}}}$). We define (see Equation (28)) the shape of $\Lambda$ as the class of the $\mathfrak{m}$}-lattice $\Lambda$ modulo scaling and modulo the maximal compact subgroup action, that is,
@@
Contrarily to the real field case, the rescaling process is much harder when the absolute value is discrete, hence the above restriction on the covolumes. Let $V_{\Lambda}$$^{\perp}$ be the subspace of the dual space $(K_{\nu}^{\;\mathbf{d}})^{*}$ of $K_{\nu}^{\;\mathbf{d}}$ consisting in the $K_{\nu}$-linear forms on $K_{\nu}^{\;\mathbf{d}}$ vanishing on $V_{\Lambda}$ and let $R_{\nu}^{\;\mathbf{d},*}$ be the standard $R_{\nu}$-lattice in $(K_{\nu}^{\;\mathbf{d}})^{*}$ generated by the dual basis of the canonical basis of $K_{\nu}^{\;\mathbf{d}}$. We also define the orthogonal $R_{\nu}$-lattice $\Lambda^{\perp}$ of the primitive $\mathfrak{m}$}-lattice $\Lambda$ by $\Lambda^{\perp}={V_{\Lambda}}^{\perp}\cap R_{\nu}^{\;\mathbf{d},*}$, which is a primitive $\mathfrak{n}$}-lattice in $(K_{\nu}^{\;\mathbf{d}})^{*}$, see Subsection 2.3. The following joint equidistribution result of the pairs of shapes of primitive $\mathfrak{m}$}-lattices and their orthogonal $\mathfrak{n}$}-lattices as their covolume tends to infinity, in the product of the quotients of the Bruhat-Tits buildings ${\cal I}_{\nu,{\mathfrak{m}}}$ and ${\cal I}_{\nu,{\mathfrak{n}}}$ by their modular groups ${\widetilde{\Gamma}}_{\mathfrak{m}}$ and ${\widetilde{\Gamma}}_{\mathfrak{n}}$, is a corollary of Theorem 1.2, see the end of Subsection 4.3 for its proof.
@@ 2.5 Refined LU decomposition by blocks
is a finite disjoint union of ${\cal U}^{\bullet}_{G}$ and finitely many left (or right) translates of ${\cal U}_{G}$. Let ${\cal S}^{\pm}_{\mathbf{d}}$ be the subgroup of $G$ consisting in the elements of $G$ that act by a permutation and a possible change of sign on the elements of the canonical basis of $K_{\nu}^{\;\mathbf{d}}$ (in order for their determinant to be $1$). Multiplying on the left an element $g\in G$ by an element in ${\cal S}^{\pm}_{\mathbf{d}}$ amounts to permuting the rows of $g$ by the inverse of the associated permutation and possibly changing their sign. Since the rank of the submatrix consisting in the first $\mathfrak{m}$} columns of any invertible ${\mathfrak{m}}\times{\mathfrak{m}}$ matrix is equal to $\mathfrak{m}$}, we have ${\cal U}^{\bullet}_{G}\subset{\cal S}^{\pm}_{\mathbf{d}}(G\smallsetminus{\cal U}^{\bullet}_{G})$. Hence up to left translations by the finitely many elements in $Z^{\prime}$ or ${\cal S}^{\pm}_{\mathbf{d}}$, in order to study equidistribution properties of natural families of points in (products of) homogeneous spaces of $G$, we may concentrate on the study of these points in the part corresponding to ${\cal U}_{G}$.
@@ 2.6 Refined LU decomposition by blocks and partial lattices
defined as follows. Let $\Lambda\in\operatorname{Lat}_{k}(V)$ with $\overline{\operatorname{Covol}}(\Lambda)\in q_{v}^{\;k{\mathbb{Z}}}$. By Lemma 2.7, choose an element $g\in\operatorname{GL}(V_{{\cal O}_{\nu}})$ such that we have $gV_{\Lambda}=K_{\nu}f_{1}+\dots+K_{\nu}f_{k}$. Note that $g$ preserves the covolume of partial $R_{\nu}$-lattices, since it maps the unit ball $V_{\Lambda}\cap V_{{\cal O}_{\nu}}$ of $V_{\Lambda}$ to the unit ball ${\cal O}_{\nu}f_{1}+\dots+{\cal O}_{\nu}f_{k}$ of $K_{\nu}f_{1}+\dots+K_{\nu}f_{k}$. Let $\Theta:K_{\nu}f_{1}+\dots+K_{\nu}f_{k}\rightarrow K_{\nu}^{\;k}$ be the isometric (hence Haar measure preserving) $K_{\nu}$-linear isomorphism mapping $(f_{1},\dots,f_{k})$ to the canonical basis of $K_{\nu}^{\;k}$, that preserves the covolume of full $R_{\nu}$-lattices. We define
@@ 3.3 The space of correlated pairs of lattices
For every $k\in{\mathbb{N}}\smallsetminus\{0\}$, let $(e_{1},\dots,e_{k})$ be the canonical $K_{\nu}$-basis of $K_{\nu}^{\;k}$. Let $\Lambda\in\operatorname{Lat}^{1}_{k}$. For every $R_{\nu}$-basis $(b_{1},\dots,b_{k})$ of $\Lambda$, which is also a $K_{\nu}$-basis of $K_{\nu}^{\;k}$, let $g\in\operatorname{GL}_{k}(R_{v})$ be the unique $K_{\nu}$-linear map sending $(e_{1},\dots,e_{k})$ to $(b_{1},\dots,b_{k})$. Note that $\det g\in{\cal O}_{\nu}^{\;\times}$ since $|\det g|=\overline{\operatorname{Covol}}(\Lambda)=1$. The image $(\det g)\,{\mathbb{F}}_{q}^{\;\times}$ of $\det g$ by the canonical projection ${\cal O}_{\nu}^{\;\times}\rightarrow{\cal O}_{\nu}^{\;\times}/{\mathbb{F}}_{q}^{\;\times}$ does not depend on the choice of the $R_{\nu}$-basis of $\Lambda$, since the change of $R_{\nu}$-basis matrix belongs to $\operatorname{GL}_{k}(R_{\nu})$, hence has determinant in $R_{\nu}^{\;\times}={\mathbb{F}}_{q}^{\;\times}$. We define
@@
Note that the ratio $\frac{a}{a^{\prime}}$ of two elements $a=\lambda{\mathbb{F}}_{q}^{\;\times}$ and $a^{\prime}=\lambda^{\prime}{\mathbb{F}}_{q}^{\;\times}$ of $K_{\nu}^{\;\times}/{\mathbb{F}}_{q}^{\;\times}$ is a well-defined element $\frac{a}{a^{\prime}}=\frac{\lambda}{\lambda^{\prime}}\,{\mathbb{F}}_{q}^{\;\times}$ of $K_{\nu}^{\;\times}/\,{\mathbb{F}}_{q}^{\;\times}$, and we will as usual also denote by ${\mathbb{F}}_{q}^{\;\times}$ the class of $1$ in $K_{\nu}^{\;\times}/\,{\mathbb{F}}_{q}^{\;\times}$. Let us define
@@
Since the stabilizer of $(R_{\nu}^{\;{\mathfrak{m}}},R_{\nu}^{\;{\mathfrak{n}}})$ by the action of $G^{\prime\prime}$ defined in Equation (43) is exactly the discrete subgroup $G^{\prime\prime}(R_{\nu})$, and since $\pi_{{\mathfrak{m}},{\mathfrak{n}}}$ is the orbital map of the above action at $(R_{\nu}^{\;{\mathfrak{m}}},R_{\nu}^{\;{\mathfrak{n}}})$, we will identify from now on the quotient space $G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ and the subspace $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$ by the homeomorphism $g\,G^{\prime\prime}(R_{\nu})\mapsto\pi_{{\mathfrak{m}},{\mathfrak{n}}}(g)=g(R_{\nu}^{\;{\mathfrak{m}}},R_{\nu}^{\;{\mathfrak{n}}})$.
Recall that $G^{\prime\prime}$ is endowed with its Haar measure $\mu_{G^{\prime\prime}}$ normalized so that $\mu_{G^{\prime\prime}}(G^{\prime\prime}({\cal O}_{\nu}))=1$. As said at the beginning of Subsection 3.2, we endow $G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ with the unique $G^{\prime\prime}$-invariant measure such that the covering map $G^{\prime\prime}\rightarrow G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ locally preserves the measure. We denote by $\mu_{\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}}$ the corresponding measure on $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$.
@@ 4.1 The correspondence between primitive partial lattices and integral group elements
Let $\Gamma=\operatorname{SL}_{\mathbf{d}}(R_{\nu})$ be the modular group of integral points of $G=\operatorname{SL}_{\mathbf{d}}(K_{\nu})$. The aim of this subsection is to naturally and injectively associate elements in $\Gamma$ to primitive $\mathfrak{m}$}-lattices in $R_{\nu}^{\;\mathbf{d}}$. 11todo: 1check if need introduce left cosets of $\Gamma$ for primitive lattices with wrong congruence of covolume, as GN still works We start by introducing the subsets of the group $G$ and of the moduli spaces $\operatorname{{\cal P}\!{\cal L}}_{{\mathfrak{m}},\mathbf{d}}$, $\operatorname{Gr}_{{\mathfrak{m}},\mathbf{d}}$, $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$ which will be technically useful.
@@
By Equation (52) and by the measure preserving indentification of $G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ and $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$ done in Subsection 3.3 (just before Lemma 3.3), we have

### v3

@@ 1 Introduction
We endow $K_{\nu}^{\;\mathbf{d}}$ with the supremum norm, and any $K_{\nu}$-linear subspace of $K_{\nu}^{\;\mathbf{d}}$ with its induced norm and with its associated normalized Haar measure (giving mass one to its closed unit ball), see Subsection 2.2. A partial $R_{\nu}$-lattice $\Lambda$ of rank $\mathfrak{m}$} (or $\mathfrak{m}$}-lattice for short) in $K_{\nu}^{\;\mathbf{d}}$ is a discrete free $R_{\nu}$-submodule of rank $\mathfrak{m}$} generating an $\mathfrak{m}$}-dimensional $K_{\nu}$-vector subspace $V_{\Lambda}$ of $K_{\nu}^{\;\mathbf{d}}$. We denote by $\operatorname{Covol}(\Lambda)$ the covolume of $\Lambda$ in $V_{\Lambda}$. We say that $\Lambda$ is primitive if it is a free direct factor of $R_{\nu}^{\;\mathbf{d}}$. Among all the definitions of primitiveness that were equivalent for partial $\mathbb{Z}$}-lattices in ${\mathbb{R}}^{\mathbf{d}}$ and non longer are, this turns out to be the appropriate one. We denote by $\operatorname{{\cal P}\!{\cal L}}_{{\mathfrak{m}},\mathbf{d}}$ the set of primitive $\mathfrak{m}$}-lattices in $K_{\nu}^{\;\mathbf{d}}$. For every point $y$ of every measurable space, we denote by $\Delta_{y}$ the unit Dirac mass at $y$.
@@
Let $\Lambda\in\operatorname{{\cal P}\!{\cal L}}_{{\mathfrak{m}},\mathbf{d}}$ be such that there exists $i\in{\mathbb{Z}}$ with $\frac{\operatorname{Covol}(\Lambda)}{\operatorname{Covol}(R_{\nu}^{\;{\mathfrak{m}}})}=q_{\nu}^{\;\operatorname{lcm}({\mathfrak{m}},{\mathfrak{n}})\,i}$ and let $g\in\operatorname{GL}_{\mathbf{d}}({\cal O}_{\nu})$ sending $V_{\Lambda}$ to $K_{\nu}^{\;{\mathfrak{m}}}\times\{0\}$ (so that $g\Lambda$ becomes a full $R_{\nu}$-lattice of $K_{\nu}^{\;{\mathfrak{m}}}$). We define (see Equation (28)) the shape of $\Lambda$ as the class of the $\mathfrak{m}$}-lattice $\Lambda$ modulo scaling and modulo the action of the maximal compact subgroup $\operatorname{GL}_{\mathbf{d}}({\cal O}_{\nu})$, that is,
@@
Contrarily to the real field case, the rescaling process is much harder when the absolute value is discrete, hence the above restriction on the covolumes. Let $V_{\Lambda}$$^{\perp}$ be the subspace of the dual space $(K_{\nu}^{\;\mathbf{d}})^{*}$ of $K_{\nu}^{\;\mathbf{d}}$ consisting in the $K_{\nu}$-linear forms on $K_{\nu}^{\;\mathbf{d}}$ vanishing on $V_{\Lambda}$ and let $R_{\nu}^{\;\mathbf{d},*}$ be the standard $R_{\nu}$-lattice in $(K_{\nu}^{\;\mathbf{d}})^{*}$ generated as an $R_{\nu}$-module by the dual basis of the canonical basis of $K_{\nu}^{\;\mathbf{d}}$. We define the orthogonal $R_{\nu}$-lattice $\Lambda^{\perp}$ of the primitive $\mathfrak{m}$}-lattice $\Lambda$ by $\Lambda^{\perp}={V_{\Lambda}}^{\perp}\cap R_{\nu}^{\;\mathbf{d},*}$, which is a primitive $\mathfrak{n}$}-lattice in $(K_{\nu}^{\;\mathbf{d}})^{*}$, see Subsection 2.3. The following joint equidistribution result of the pairs of shapes of primitive $\mathfrak{m}$}-lattices and their orthogonal $\mathfrak{n}$}-lattices as their covolume tends to infinity, in the product of the quotients of the Bruhat-Tits buildings ${\cal I}_{\nu,{\mathfrak{m}}}$ and ${\cal I}_{\nu,{\mathfrak{n}}}$ by their modular groups ${\widetilde{\Gamma}}_{\mathfrak{m}}$ and ${\widetilde{\Gamma}}_{\mathfrak{n}}$, is a corollary of Theorem 1.2. See the end of Subsection 4.3 for its proof.
@@ 2.5 Refined LU decomposition by blocks
is a finite disjoint union of ${\cal U}^{\bullet}_{G}$ and finitely many left (or right) translates of ${\cal U}_{G}$. Let ${\cal S}^{\pm}_{\mathbf{d}}$ be the subgroup of $G$ consisting in the elements of $G$ that act by a permutation and a possible change of sign on the elements of the canonical basis of $K_{\nu}^{\;\mathbf{d}}$ (in order for their determinant to be $1$). Multiplying on the left an element $g\in G$ by an element in ${\cal S}^{\pm}_{\mathbf{d}}$ amounts to permuting the rows of $g$ by the inverse of the associated permutation and possibly changing their sign. Since the rank of the submatrix consisting in the first $\mathfrak{m}$} columns of any invertible $\mathbf{d}\times\mathbf{d}$ matrix is equal to $\mathfrak{m}$}, we have ${\cal U}^{\bullet}_{G}\subset{\cal S}^{\pm}_{\mathbf{d}}(G\smallsetminus{\cal U}^{\bullet}_{G})$. Hence up to left translations by the finitely many elements in $Z^{\prime}$ or ${\cal S}^{\pm}_{\mathbf{d}}$, in order to study equidistribution properties of natural families of points in (products of) homogeneous spaces of $G$, we may concentrate on the study of these points in the part corresponding to ${\cal U}_{G}$.
@@ 2.6 Refined LU decomposition by blocks and partial lattices
defined as follows. Let $\Lambda\in\operatorname{Lat}_{k}(V)$ with $\overline{\operatorname{Covol}}(\Lambda)\in q_{\nu}^{\;k{\mathbb{Z}}}$. By Lemma 2.7, choose an element $g\in\operatorname{GL}(V_{{\cal O}_{\nu}})$ such that we have $gV_{\Lambda}=K_{\nu}f_{1}+\dots+K_{\nu}f_{k}$. Note that $g$ preserves the covolume of partial $R_{\nu}$-lattices, since it maps the unit ball $V_{\Lambda}\cap V_{{\cal O}_{\nu}}$ of $V_{\Lambda}$ to the unit ball ${\cal O}_{\nu}f_{1}+\dots+{\cal O}_{\nu}f_{k}$ of $K_{\nu}f_{1}+\dots+K_{\nu}f_{k}$. Let $\Theta:K_{\nu}f_{1}+\dots+K_{\nu}f_{k}\rightarrow K_{\nu}^{\;k}$ be the isometric (hence Haar measure preserving) $K_{\nu}$-linear isomorphism mapping $(f_{1},\dots,f_{k})$ to the canonical basis of $K_{\nu}^{\;k}$, that preserves the covolume of full $R_{\nu}$-lattices. We define
@@ 3.3 The space of correlated pairs of lattices
For every $k\in{\mathbb{N}}\smallsetminus\{0\}$, let $(e_{1},\dots,e_{k})$ be the canonical $K_{\nu}$-basis of $K_{\nu}^{\;k}$. Let $\Lambda\in\operatorname{Lat}^{1}_{k}$. For every $R_{\nu}$-basis $(b_{1},\dots,b_{k})$ of $\Lambda$, which is also a $K_{\nu}$-basis of $K_{\nu}^{\;k}$, let $g\in\operatorname{GL}_{k}(K_{\nu})$ be the unique $K_{\nu}$-linear map sending $(e_{1},\dots,e_{k})$ to $(b_{1},\dots,b_{k})$. Note that $\det g\in{\cal O}_{\nu}^{\;\times}$ since $|\det g|=\overline{\operatorname{Covol}}(\Lambda)=1$. The image $(\det g)\,{\mathbb{F}}_{q}^{\;\times}$ of $\det g$ by the canonical projection ${\cal O}_{\nu}^{\;\times}\rightarrow{\cal O}_{\nu}^{\;\times}/{\mathbb{F}}_{q}^{\;\times}$ does not depend on the choice of the $R_{\nu}$-basis of $\Lambda$, since the change of $R_{\nu}$-basis matrix belongs to $\operatorname{GL}_{k}(R_{\nu})$, hence has determinant in $R_{\nu}^{\;\times}={\mathbb{F}}_{q}^{\;\times}$. We define
@@
Note that the ratio $\frac{a}{a^{\prime}}$ of two elements $a=\lambda{\mathbb{F}}_{q}^{\;\times}$ and $a^{\prime}=\lambda^{\prime}{\mathbb{F}}_{q}^{\;\times}$ of ${\cal O}_{\nu}^{\;\times}/{\mathbb{F}}_{q}^{\;\times}$ is a well-defined element $\frac{a}{a^{\prime}}=\frac{\lambda}{\lambda^{\prime}}\,{\mathbb{F}}_{q}^{\;\times}$ of ${\cal O}_{\nu}^{\;\times}/\,{\mathbb{F}}_{q}^{\;\times}$, and we will as usual also denote by ${\mathbb{F}}_{q}^{\;\times}$ the class of $1$ in ${\cal O}_{\nu}^{\;\times}/\,{\mathbb{F}}_{q}^{\;\times}$. Let us define
@@
Since the stabilizer of $(R_{\nu}^{\;{\mathfrak{m}}},R_{\nu}^{\;{\mathfrak{n}}})$ by the action of $G^{\prime\prime}$ defined in Equation (43) is exactly the discrete subgroup $G^{\prime\prime}(R_{\nu})$, and since $\pi_{{\mathfrak{m}},{\mathfrak{n}}}$ is the orbital map of the above action at $(R_{\nu}^{\;{\mathfrak{m}}},R_{\nu}^{\;{\mathfrak{n}}})$, we will identify from now on the quotient space $G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ and the subspace $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$ by the $G^{\prime\prime}$-equivariant homeomorphism $g\,G^{\prime\prime}(R_{\nu})\mapsto\pi_{{\mathfrak{m}},{\mathfrak{n}}}(g)=g(R_{\nu}^{\;{\mathfrak{m}}},R_{\nu}^{\;{\mathfrak{n}}})$.
Recall that $G^{\prime\prime}$ is endowed with its Haar measure $\mu_{G^{\prime\prime}}$ normalized so that we have $\mu_{G^{\prime\prime}}(G^{\prime\prime}({\cal O}_{\nu}))=1$. As said at the beginning of Subsection 3.2, we endow $G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ with the unique $G^{\prime\prime}$-invariant measure such that the covering map $G^{\prime\prime}\rightarrow G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ locally preserves the measure. We denote by $\mu_{\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}}$ the corresponding measure on $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$, which is $G^{\prime\prime}$-invariant.
@@ 4.1 The correspondence between primitive partial lattices and integral group elements
Let $\Gamma=\operatorname{SL}_{\mathbf{d}}(R_{\nu})$ be the modular group of integral points of $G=\operatorname{SL}_{\mathbf{d}}(K_{\nu})$. The aim of this subsection is to naturally and injectively associate elements in $\Gamma$ to primitive $\mathfrak{m}$}-lattices in $R_{\nu}^{\;\mathbf{d}}$. We start by introducing the subsets of the group $G$ and of the moduli spaces $\operatorname{{\cal P}\!{\cal L}}_{{\mathfrak{m}},\mathbf{d}}$, $\operatorname{Gr}_{{\mathfrak{m}},\mathbf{d}}$, $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$ which will be technically useful.
@@
By Equation (52) and by the measure preserving identification of $G^{\prime\prime}/G^{\prime\prime}(R_{\nu})$ and $\operatorname{Lat}^{1}_{{\mathfrak{m}},{\mathfrak{n}}}$ done in Subsection 3.3 (just before Lemma 3.3), we have